OpenStudy (anonymous):

Part 1: If you were solving a system of equations and you came to a statement like 1 = 3, what do you know about the solution(s) to this system? Part 2: Solve the following system and show all of your work. x − 2y = 14 x + 3y = 9

Parth (parthkohli):

1) You will come to a conclusion that there's no solution to the system :)

Parth (parthkohli):

2) You can solve by elimination.

Parth (parthkohli):

Subtract the equations and solve for \(y\), then solve for \(x\).

OpenStudy (anonymous):

@ParthKohli do you mind helping me solve by elimination? I don't know how to do these

Parth (parthkohli):

Subtract both equations: \( \color{Black}{\Rightarrow -5y = 5}\) \( \color{Black}{\Rightarrow y = -1}\) Now as you have solved for \(y\), you can solve for \(x\) by substituting y in any equation.

Parth (parthkohli):

\( \color{Black}{\Rightarrow x + 3(-1) = 9 }\) Can you manage now?

OpenStudy (anonymous):

I don't know what to do from here...

Parth (parthkohli):

\( \color{Black}{\Rightarrow x + (-3) = 9 }\) \( \color{Black}{\Rightarrow x - 3 = 9 }\) Now?

OpenStudy (anonymous):

No, I don't know the steps or process. Plus, I have to get to 14 in the first equation?

Parth (parthkohli):

No need.

OpenStudy (anonymous):

? hold on i'll show you what I put

OpenStudy (anonymous):